After deciding to buy a new car, you can either lease the car or purchase it on a 3-year loan. The car you wish to buy costs $43,000. The dealer has a special leasing arrangement where you pay $4,300 today and $505 per month for the next 3 years. If you purchase the car, you will pay it off in monthly payments over the next 3 years at an APR of 6%. You believe you will be able to sell the car for $28,000 in 3 years. Should you lease the car? What break-even resale price in 3 years should make you indifferent between buying and leasing?
Answer #1


solution: monthlly rate = 0.06/12 number of period = 3*12 = 36 Present value of Purchasing car: Purchase cost Less : Present